OpenStudy (anonymous):

The sides of a square are three to the power of two sevenths inches long. What is the area of the square? nine to the power of four sevenths square inches three to the power of four sevenths square inches three to the power of the fraction four over forty nine square inches nine to the power of the fraction four over forty nine square inches

OpenStudy (anonymous):

I will give a medal if anyone can help me

OpenStudy (anonymous):

jigglypuff314

OpenStudy (anonymous):

#ParthKohli

jhonyy9 (jhonyy9):

area of a square = side time side - yes ?

jhonyy9 (jhonyy9):

the length of a side is 3^(3/7) - so i dont know than i have understood it your text about this side length right - or hence is right sure ?

jhonyy9 (jhonyy9):

sorry 3^(2/7) - not (3/7) - my mistaque

jhonyy9 (jhonyy9):

so than 3^(2/7) *3^(2/7) = 3^(2/7 +2/7) = 3^((2+2)/7) = 3^(4/7) hope this will help you yes ?

OpenStudy (anonymous):

yes it did thank you so much

jhonyy9 (jhonyy9):

np was my pleasure god luck and so anytime bye